# The Ginsburg–Sands theorem and computability theory

### v1

@@ Abstract.
The Ginsburg–Sands theorem from topology states that every infinite topological space has an infinite subspace homeomorphic to exactly one of the following five topologies on $\omega$: indiscrete, discrete, initial segment, final segment, and cofinite. The original proof is nonconstructive, and features an interesting application of Ramsey’s theorem for pairs ($\mathsf{RT}^{2}_{2}$). We analyze this principle in computability theory and reverse mathematics, using Dorais’s formalization of CSC spaces. Among our results are that the Ginsburg-Sands theorem for CSC spaces is equivalent to $\mathsf{ACA}_{0}$, while for Hausdorff spaces it is provable in $\mathsf{RCA}_{0}$. Furthermore, if we enrich a CSC space by adding the closure operator on points, then the Ginsburg-Sands theorem turns out to be equivalent to the chain/antichain principle ($\mathsf{CAC}$). The most surprising case is that of the Ginsburg-Sands theorem restricted to $T_{1}$ spaces. Here, we show that the principle lies strictly between $\mathsf{ACA}_{0}$ and $\mathsf{RT}^{2}_{2}$, yielding arguably the first natural theorem from outside logic to occupy this interval. As part of our analysis of the $T_{1}$ case we introduce a new class of purely combinatorial principles below $\mathsf{ACA}_{0}$ and not implied by $\mathsf{RT}^{2}_{2}$ which form a strict hierarchy generalizing the stable Ramsey’s theorem for pairs ($\mathsf{SRT}^{2}_{2}$). We show that one of these, the $\Sigma^{0}_{2}$ subset principle ($\Sigma^{0}_{2}\text{-}\mathsf{Subset}$), has the property that it, together with the cohesive principle ($\mathsf{COH}$), is equivalent over $\mathsf{RCA}_{0}$ to the Ginsburg–Sands theorem for $T_{1}$ CSC spaces.
The authors were partially supported by a Focused Research Group grant from the National Science Foundation of the United States, DMS-1854355. They thank Natasha Dobrinen, Denis Hirschfeldt, Ludovic Levy Patey, and Keita Yokoyama for insightful discussions during the preparation of this article.
@@ 1. Introduction
In their 1979 paper “Minimal infinite topological spaces” [17], Ginsburg and Sands proved the following classification theorem for topological spaces, described by Marron and McMaster [29, pp. 26–27] as a “[significant] application of Ramsey’s theorem”.
@@
Since the result is not especially well-known, we include the original proof from [17] in Section 2 below. As Ginsburg and Sands explain [17, p. 574], any infinite subspace of a topological space homeomorphic to one of the above five is itself homeomorphic to the same, and no two distinct such spaces can be homeomorphic to each other. Thus, the above topologies are “minimal” in the same sense that a homogeneous set is for a given coloring of tuples of integers: if a homogeneous set has color $i$ for some coloring then so does every subset, and no homogeneous set can simultaneously have colors $i$ and $j\neq i$.
@@
Computability theory is concerned with the question of which subsets of the natural numbers can be specified by an algorithm, and by extension, seeks to determine the algorithmic properties of mathematical theorems that can be phrased in terms of, or represented by, countable objects. Reverse mathematics, on the other hand, investigates the foundational principles underlying mathematical theorems by identifying the minimal logical axioms necessary for their proof, commonly using subsystems of second-order arithmetic as benchmarks of logical strength. There is a well-understood relationship between these two endeavors, with notions and results in one often leading to notions and results in the other. This combined perspective often provides deeper insights into the logical underpinnings of mathematical theorems, particularly with respect to the combinatorial tools and methods needed to prove them (see, e.g., [19], p. 65).
@@
Ramsey’s theorem, and many related principles from combinatorics, model theory, and set theory have been studied extensively in this context. A partial survey appears in [15, Chapters 8 and 9]. Topological theorems of various kinds have also been considered, but because of the need for countable representations a lot here depends on how exactly topological spaces are coded. Thus, different approaches exist for studying metric spaces (see [15], Sections 10.1–10.5 for an overview), certain uncountable spaces (see Mummert [35] and Mummert and Lempp [26]), and countable spaces with countably based topologies (see Dorais [10]). Topology has also been studied using higher-order reverse mathematics (Hunter [23], Normann and Sanders [36], and Sanders [40, 41]). A more detailed discussion of these various viewpoints appears in [15, Section 10.8].
Our investigation of the Ginsburg–Sands theorem will use Dorais’s formalization of countable second-countable (CSC) spaces in classical (second-order) reverse mathematics. The original account [10] of this is unpublished, so we review some of the main definitions and the features most relevant for our purposes in Section 3. Additional background can be found in [15, Section 10.8.1]. Other recent applications besides our own include the work of Shafer [43] and DeLapo [9].
Our goal is to better understand the Ginsburg–Sands theorem in isolation, but also with respect to Ramsey’s theorem for pairs, which figures in its proof. For the remainder of this section, then, we recall the statement of Ramsey’s theorem, and of some of its consequences. Throughout, our notation will be standard unless otherwise noted. We assume familiarity with the usual benchmark subsystems $\mathsf{RCA}_{0}$, $\mathsf{WKL}_{0}$, and $\mathsf{ACA}_{0}$ of second-order arithmetic. We follow the convention of formulating all definitions in $\mathsf{RCA}_{0}$ for generality, thereby using $\mathbb{N}$ in place of $\omega$ so as to refer to the set of natural numbers in a possibly nonstandard model. For $n\geq 1$, we use $\mathsf{I}\Sigma^{0}_{n}$ to refer to the $\Sigma^{0}_{n}$ induction scheme, which consists of all formulas of the form
@@
It is easy to see that for each $n$, the principle $\mathsf{RT}^{n}_{1}$ is provable in $\mathsf{RCA}_{0}$. Hirst [21, Theorem 6.4] showed that $(\forall m)[\mathsf{RT}^{1}_{m}]$ is equivalent to $\mathsf{B}\Sigma^{0}_{2}$, the $\Sigma^{0}_{2}$ bounding scheme. We recall that if $\Gamma$ is any collection of formulas of $\mathsf{L}_{2}$, the scheme $\mathsf{B}\Gamma$ consists of all formulas of the form
@@
For any fixed, standard $m\geq 2$, the usual collapsing argument shows that $\mathsf{RT}^{n}_{m}$ is equivalent over $\mathsf{RCA}_{0}$ to $\mathsf{RT}^{n}_{2}$. By a result of Jockusch [25, Lemma 5.9], $\mathsf{RT}^{3}_{2}$ is equivalent over $\mathsf{RCA}_{0}$ to $\mathsf{ACA}_{0}$. By contrast, Seetapun (see [42], Theorem 2.1) showed that $\mathsf{RT}^{2}_{2}$ is strictly weaker than $\mathsf{ACA}_{0}$, while Liu [28, Corollary 1.6] showed that it does not imply $\mathsf{WKL}_{0}$. A modern proof of Seetapun’s theorem uses the so-called strong cone avoidance of $\mathsf{RT}^{1}_{2}$ of Dzhafarov and Jockusch [14, Lemma 3.2] (see also [15], Section 8.5.1).
@@
We mention, for completeness, that Anglès d’Auriac et al. [8, Theorem 4.15] established the fact that $\mathsf{MTT}^{2}_{2}$ does not imply $\mathsf{ACA}_{0}$. Dzhafarov and Patey [16, Theorem 4.1] showed that $\mathsf{TT}^{2}_{2}$ does not imply $\mathsf{ACA}_{0}$, which Chong et al. [4, Theorem 1.1] improved to show that $\mathsf{TT}^{2}_{2}$ does not imply even $\mathsf{WKL}_{0}$. Patey [38, Theorem 1.3] proved that $\mathsf{RT}^{2}_{2}$ does not imply $\mathsf{TT}^{2}_{2}$. It remains open whether $\mathsf{MTT}^{2}_{2}$ implies $\mathsf{WKL}_{0}$, or whether $\mathsf{TT}^{2}_{2}$ implies $\mathsf{MTT}^{2}_{2}$.
@@
Given a set $X$, a coloring $c:[X]^{2}\to 2$ is stable if $(\forall x)[\lim_{y}c(x,y)]$ exists. A set $L\subseteq X$ is limit homogeneous for $c$ if $(\exists i<2)(\forall x\in L)[\lim_{y}c(x,y)=i]$.
@@
An infinite homogeneous set for a stable coloring is trivially limit homogeneous. Conversely, an infinite limit homogeneous set can be easily thinned to an infinite homogeneous set, but proving that this can be done with only $\Sigma^{0}_{1}$ induction in $\mathsf{RCA}_{0}$ is delicate, and is a theorem of Chong, Lempp, and Yang [3, Theorem 1.4]. Thus, $\mathsf{RCA}_{0}$ proves $\mathsf{SRT}^{2}_{2}\leftrightarrow\mathsf{D}^{2}_{2}$. An alternative formulation of $\mathsf{D}^{2}_{2}$ is as the scheme consisting of all formulas of the form
@@
In [2, Theorem 9.1], it was shown that $\mathsf{COH}$ does not imply $\mathsf{SRT}^{2}_{2}$, but whether or not $\mathsf{SRT}^{2}_{2}$ implies $\mathsf{COH}$ (or equivalently, $\mathsf{RT}^{2}_{2}$) was a major open question in reverse mathematics for many years. It was finally answered in 2014 by Chong, Slaman, and Yang [5, Theorem 2.2], who constructed a nonstandard model of $\mathsf{RCA}_{0}+\mathsf{SRT}^{2}_{2}$ in which $\mathsf{COH}$ does not hold. More recently, Monin and Patey [34, Theorem 1.4], improved this to a separation via an $\omega$-model. See [15, Section 8.8] for more on the history of this problem.
@@
It is easy to see that $\mathsf{RT}^{2}_{2}$ implies $\mathsf{CAC}$, and that $\mathsf{CAC}$ implies $\mathsf{ADS}$. Hirschfeldt and Shore [20, Corollaries 2.16 and 3.12] exhibited an $\omega$-model satisfying $\mathsf{RCA}_{0}+\mathsf{CAC}$ but not $\mathsf{SRT}^{2}_{2}$, as well as an $\omega$-model satisfying $\mathsf{WKL}_{0}$ but not $\mathsf{ADS}$. They also showed ([20], Proposition 2.10) that $\mathsf{ADS}$ implies $\mathsf{COH}$. This left open whether or not $\mathsf{ADS}$ implies $\mathsf{CAC}$ over $\mathsf{RCA}_{0}$, which was finally answered, in the negative and via an $\omega$-model, by Lerman, Solomon, and Towsner [27, Theorem 1.7].
It is worth noting that $\mathsf{CAC}$ and $\mathsf{ADS}$ were also identified by Ginsburg and Sands [17, p. 575], and it is these principles, rather than $\mathsf{RT}^{2}_{2}$, that they employed in their proof of Theorem 1.1, which we present in the next section.
We follow the usual convention of also thinking of $\Pi^{1}_{2}$ principles such as those above as problems, each having a class of instances, and for each instance, a class of solutions (see [15, Chapter 3] for a general discussion). Thus, for example, the instances of $\mathsf{RT}^{2}_{2}$ are all pairs $\langle X,c\rangle$ such that $X$ is an infinite set and $c:[X]^{2}\to 2$ is a coloring, with the solutions to $c$ being all infinite homogeneous sets $H\subseteq X$.
@@ 2. A proof of the Ginsburg–Sands theorem
Fix an infinite topological space $X$. By passing to a subspace if needed, we may assume $X$ is countable. For each $x\in X$, let $\operatorname{cl}(x)$ denote the closure of $x$ in $X$, i.e., the set of all $y\in X$ such that every open set containing $y$ also contains $x$. Then, for $x,y\in\omega$, define $x\sim y$ if $\operatorname{cl}(x)=\operatorname{cl}(y)$, which is an equivalence relation on $X$. If there is an $x$ such that the equivalence class $\{y\in\omega:x\sim y\}$ is infinite, then the subspace topology on this class is an indiscrete subspace of $X$. Moving forward, we can therefore assume that all equivalence classes under $\sim$ are finite. By choosing representatives, we may further assume that $\operatorname{cl}(x)\neq\operatorname{cl}(y)$ for all $x\neq y$ in $X$, so that $X$ itself is a $T_{0}$ space. We now define a relation $\leq_{X}$ on $X$, as follows. For $x,y\in X$, set $x\leq_{X}y$ if $x\in\operatorname{cl}(y)$. Clearly, $\leq_{X}$ is reflexive and transitive, and our assumption that $X$ is $T_{0}$ implies that $\leq_{X}$ is antisymmetric. So, $\leq_{X}$ is a partial ordering. By $\mathsf{CAC}$, there exists an infinite subset of $X$ that either forms a chain under $\leq_{X}$ or an antichain under $\leq_{X}$. By regarding this as a subspace, we may assume this subset is in fact of all of $X$ for simplicity. First, suppose $X$ is a chain under $\leq_{X}$. Applying $\mathsf{ADS}$, we can fix an infinite subset of $X$ that is either an ascending or descending sequence under $\leq_{X}$. Again, by thinking of this as a subspace, we may assume this subset is in fact all of $X$. So suppose $X$ is an ascending sequence under $\leq_{X}$. Then we can find $x_{0}<x_{1}<\cdots$ in $X$ such that $x_{0}<_{X}x_{1}<_{X}\cdots$. We claim that $Y=\{x_{n}:n\in\omega\}$ under the subspace topology is homeomorphic to the final segment topology on $\omega$ via the map $x_{n}\mapsto n$. To see this, note first that if $n<m$ then every open set in $X$ containing $x_{n}$ also contains $x_{m}$ since $x_{n}\in\operatorname{cl}(x_{m})$. On the other hand, for every $n$ there is an open set containing $x_{n}$ but not any $x_{m}$ for $m<n$, since $x_{n}\notin\operatorname{cl}(x_{m})$ for any such $m$. It follows that the open sets in $Y$ are precisely the sets of the form $\{x_{m}:m\geq n\}$ for $n\in\omega$, which proves the claim. The case where $X$ is a descending sequence under $\leq_{X}$ is symmetric, and produces a subspace homeomorphic to the initial segment topology. We are thus left with the case where $X$ is an antichain. By definition of $\leq_{X}$, this means that $X$ is a $T_{1}$ space. Assume $X$ has no infinite subspace with the cofinite topology; we show it has an infinite discrete subspace. Since $X$ is $T_{1}$, every finite subset of $X$ is closed, and so every cofinite subset is open. Our assumption therefore implies that if $W$ is any infinite subspace of $X$, there exists an open set $U$ such that $U\cap W\neq\emptyset$ and $W\smallsetminus U$ is infinite. Let $U_{0}$ be any such open set for $W=X$, and let $x_{0}$ be any element of $U_{0}$. Next, suppose inductively that, for some $n\in\omega$, we have defined $x_{0},\ldots,x_{n}\in X$, along with open sets $U_{0},\ldots,U_{n}$ satisfying the following: $X\smallsetminus\bigcup_{i\leq n}U_{i}$ is infinite, and for all $i,j\leq n$, we have that $x_{i}\in U_{j}$ if and only if $i=j$. Letting $W=X\smallsetminus\bigcup_{i\leq n}U_{i}$, there exists an open set $U$ such that $U\cap W\neq\emptyset$ and $W\smallsetminus U$ is infinite. Let $x_{n+1}$ be any element of $U\cap W$, and let $U_{n+1}=U\smallsetminus\{x_{0},\ldots,x_{n}\}$. (Note that $U_{n+1}$ is open because $X$ is $T_{1}$.) Clearly, the inductive conditions are maintained. In this way, we thus define an infinite subspace $Y=\{x_{n}:n\in\omega\}$ of $X$, and this subspace is discrete since for each $n$ we have that $U_{n}\cap Y=\{x_{n}\}$ by construction. ∎

### v2

@@ Abstract.
The Ginsburg–Sands theorem from topology states that every infinite topological space has an infinite subspace homeomorphic to exactly one of the following five topologies on $\omega$: indiscrete, discrete, initial segment, final segment, and cofinite. The original proof is nonconstructive, and features an interesting application of Ramsey’s theorem for pairs ($\mathsf{RT}^{2}_{2}$). We analyze this principle in computability theory and reverse mathematics, using Dorais’s formalization of CSC spaces. Among our results are that the Ginsburg–Sands theorem for CSC spaces is equivalent to $\mathsf{ACA}_{0}$, while for Hausdorff spaces it is provable in $\mathsf{RCA}_{0}$. Furthermore, if we enrich a CSC space by adding the closure operator on points, then the Ginsburg–Sands theorem turns out to be equivalent to the chain/antichain principle ($\mathsf{CAC}$). The most surprising case is that of the Ginsburg–Sands theorem restricted to $T_{1}$ spaces. Here, we show that the principle lies strictly between $\mathsf{ACA}_{0}$ and $\mathsf{RT}^{2}_{2}$, yielding arguably the first natural theorem from outside logic to occupy this interval. As part of our analysis of the $T_{1}$ case we introduce a new class of purely combinatorial principles below $\mathsf{ACA}_{0}$ and not implied by $\mathsf{RT}^{2}_{2}$ which form a strict hierarchy generalizing the stable Ramsey’s theorem for pairs ($\mathsf{SRT}^{2}_{2}$). We show that one of these, the $\Sigma^{0}_{2}$ subset principle ($\Sigma^{0}_{2}\text{-}\mathsf{Subset}$), has the property that it, together with the cohesive principle ($\mathsf{COH}$), is equivalent over $\mathsf{RCA}_{0}$ to the Ginsburg–Sands theorem for $T_{1}$ CSC spaces.
The authors were partially supported by a Focused Research Group grant from the National Science Foundation of the United States, DMS-1854355. They thank Natasha Dobrinen, Denis Hirschfeldt, Ludovic Levy Patey, Theodore Slaman, and Keita Yokoyama for insightful discussions during the preparation of this article, and the anonymous referee for valuable comments and suggestions for improvement.
@@ 1. Introduction
In their 1979 paper “Minimal infinite topological spaces” [18], Ginsburg and Sands proved the following classification theorem for topological spaces, described by Marron and McMaster [31, pp. 26–27] as a “[significant] application of Ramsey’s theorem”.
@@
Since the result is not especially well-known, we include the original proof from [18] in Section 2 below. As Ginsburg and Sands explain [18, p. 574], any infinite subspace of a topological space homeomorphic to one of the above five is itself homeomorphic to the same, and no two distinct such spaces can be homeomorphic to each other. Thus, the above topologies are “minimal” in the same sense that a homogeneous set is for a given coloring of tuples of integers: if a homogeneous set has color $i$ for some coloring then so does every subset, and no homogeneous set can simultaneously have colors $i$ and $j\neq i$.
@@
Computability theory is concerned with the question of which subsets of the natural numbers can be specified by an algorithm, and by extension, seeks to determine the algorithmic properties of mathematical theorems that can be phrased in terms of, or represented by, countable objects. Reverse mathematics, on the other hand, seeks to classify the logical strengths of mathematical theorems by identifying the minimal axioms necessary for their proofs, commonly using subsystems of second-order arithmetic as benchmarks. There is a well-understood relationship between these two endeavors, with notions and results in one often leading to notions and results in the other. This combined perspective often provides deeper insights into the logical underpinnings of mathematical theorems, particularly with respect to the combinatorial tools and methods needed to prove them (see, e.g., [20], p. 65).
@@
Ramsey’s theorem, and many related principles from combinatorics, model theory, and set theory have been studied extensively in this context. A partial survey appears in [16, Chapters 8 and 9]. Topological theorems of various kinds have also been considered, but because of the need for countable representations a lot here depends on how exactly topological spaces are coded. Thus, different approaches exist for studying metric spaces (see [16], Sections 10.1–10.5 for an overview), certain uncountable spaces (see Mummert [38] and Mummert and Lempp [28]), and countable spaces with countably based topologies (see Dorais [10]). Topology has also been studied using higher-order reverse mathematics (Hunter [24], Normann and Sanders [39], and Sanders [43, 44]). A more detailed discussion of these various viewpoints appears in [16, Section 10.8].
Our investigation of the Ginsburg–Sands theorem will use Dorais’s formalization of countable second-countable (CSC) spaces in classical (second-order) reverse mathematics. The original account [10] of this is unpublished, so we review some of the main definitions and the features most relevant for our purposes in Section 3. Additional background can be found in [16, Section 10.8.1]. Other recent applications besides our own include the work of Shafer [46] and DeLapo [9].
Our goal is to better understand the Ginsburg–Sands theorem in isolation, but also with respect to Ramsey’s theorem for pairs, which figures in its proof. For the remainder of this section, then, we recall the statement of Ramsey’s theorem, and of some of its consequences. Throughout, our notation will be standard unless otherwise noted. We assume familiarity with the usual benchmark subsystems $\mathsf{RCA}_{0}$, $\mathsf{WKL}_{0}$, and $\mathsf{ACA}_{0}$ of second-order arithmetic. For $n\geq 1$, we use $\mathsf{I}\Sigma^{0}_{n}$ to refer to the $\Sigma^{0}_{n}$ induction scheme, which consists of all formulas of the form
@@
where $\varphi(x)$ is a $\Sigma^{0}_{n}$ formula in the language of second-order arithmetic. All implications between principles and schemes in the sequel will be assumed to be over $\mathsf{RCA}_{0}$, unless otherwise noted. We also follow the convention of formulating all definitions in $\mathsf{RCA}_{0}$ for generality, thereby using $\mathbb{N}$ in place of $\omega$ so as to refer to the set of natural numbers in a possibly nonstandard model.
@@
It is easy to see that for each $m$, the principle $\mathsf{RT}^{1}_{m}$ is provable in $\mathsf{RCA}_{0}$. Hirst [22, Theorem 6.4] showed that $(\forall m)[\mathsf{RT}^{1}_{m}]$ is equivalent (over $\mathsf{RCA}_{0}$) to $\mathsf{B}\Sigma^{0}_{2}$, the $\Sigma^{0}_{2}$ bounding scheme. We recall that if $\Gamma$ is any collection of formulas of $\mathsf{L}_{2}$, the scheme $\mathsf{B}\Gamma$ consists of all formulas of the form
@@
For any fixed, standard $m\geq 2$, the usual collapsing argument shows that $\mathsf{RT}^{n}_{m}$ is equivalent over $\mathsf{RCA}_{0}$ to $\mathsf{RT}^{n}_{2}$. By a result of Jockusch [25, Lemma 5.9], $\mathsf{RT}^{3}_{2}$ is equivalent over $\mathsf{RCA}_{0}$ to $\mathsf{ACA}_{0}$. By contrast, Seetapun (see [45], Theorem 2.1) showed that $\mathsf{RT}^{2}_{2}$ is strictly weaker than $\mathsf{ACA}_{0}$, while Liu [30, Corollary 1.6] showed that it does not imply $\mathsf{WKL}_{0}$. A modern proof of Seetapun’s theorem uses the so-called strong cone avoidance of $\mathsf{RT}^{1}_{2}$ of Dzhafarov and Jockusch [15, Lemma 3.2] (see also [16], Section 8.5.1).
@@
We mention, for completeness, that Anglès d’Auriac et al. [8, Theorem 4.15] established the fact that $\mathsf{MTT}^{2}_{2}$ does not imply $\mathsf{ACA}_{0}$. Dzhafarov and Patey [17, Theorem 4.1] showed that $\mathsf{TT}^{2}_{2}$ does not imply $\mathsf{ACA}_{0}$, which Chong et al. [4, Theorem 1.1] improved to show that $\mathsf{TT}^{2}_{2}$ does not imply even $\mathsf{WKL}_{0}$. Patey [41, Theorem 1.3] proved that $\mathsf{RT}^{2}_{2}$ does not imply $\mathsf{TT}^{2}_{2}$. It remains open whether $\mathsf{MTT}^{2}_{2}$ implies $\mathsf{WKL}_{0}$, or whether $\mathsf{TT}^{2}_{2}$ implies $\mathsf{MTT}^{2}_{2}$.
@@
Given a set $X$, a coloring $c:[X]^{2}\to 2$ is stable if $(\forall x)[\lim_{y}c(x,y)$ exists]]]]. A set $L\subseteq X$ is limit homogeneous for $c$ if $(\exists i<2)(\forall x\in L)[\lim_{y}c(x,y)=i]$.
@@
An infinite homogeneous set for a stable coloring is trivially limit homogeneous. Conversely, an infinite limit homogeneous set can be easily thinned to an infinite homogeneous set, but this thinning process in general requires $\mathsf{B}\Sigma^{0}_{2}$ (see Dzhafarov, Hirschfeldt, and Reitzes [13, Definition 6.1 and Proposition 6.2]). However, Chong, Lempp, and Yang [3, Theorem 1.4] showed that $\mathsf{D}^{2}_{2}$ implies $\mathsf{B}\Sigma^{0}_{2}$, and hence that $\mathsf{RCA}_{0}$ proves $\mathsf{SRT}^{2}_{2}\leftrightarrow\mathsf{D}^{2}_{2}$. An alternative formulation of $\mathsf{D}^{2}_{2}$ is as the scheme consisting of all formulas of the form
@@
In [2, Theorem 9.1], it was shown that $\mathsf{COH}$ does not imply $\mathsf{SRT}^{2}_{2}$, but whether or not $\mathsf{SRT}^{2}_{2}$ implies $\mathsf{COH}$ (or equivalently, $\mathsf{RT}^{2}_{2}$) was a major open question in reverse mathematics for many years. It was finally answered in 2014 by Chong, Slaman, and Yang [5, Theorem 2.2], who constructed a nonstandard model of $\mathsf{RCA}_{0}+\mathsf{SRT}^{2}_{2}$ in which $\mathsf{COH}$ does not hold. More recently, Monin and Patey [36, Theorem 1.4], improved this to a separation via an $\omega$-model. See [16, Section 8.8] for more on the history of this problem.
@@
It is easy to see that $\mathsf{RT}^{2}_{2}$ implies $\mathsf{CAC}$, and that $\mathsf{CAC}$ implies $\mathsf{ADS}$. Hirschfeldt and Shore [21, Corollaries 2.16 and 3.12] exhibited an $\omega$-model satisfying $\mathsf{RCA}_{0}+\mathsf{CAC}$ but not $\mathsf{SRT}^{2}_{2}$, as well as an $\omega$-model satisfying $\mathsf{WKL}_{0}$ but not $\mathsf{ADS}$. They also showed ([21], Proposition 2.10) that $\mathsf{ADS}$ implies $\mathsf{COH}$. This left open whether or not $\mathsf{ADS}$ implies $\mathsf{CAC}$ over $\mathsf{RCA}_{0}$, which was finally answered, in the negative and via an $\omega$-model, by Lerman, Solomon, and Towsner [29, Theorem 1.7].
It is worth noting that $\mathsf{CAC}$ and $\mathsf{ADS}$ were also identified by Ginsburg and Sands [18, p. 575], and it is these principles, rather than $\mathsf{RT}^{2}_{2}$, that they employed in their proof of Theorem 1.1, which we present in the next section.
We follow the usual convention of also thinking of $\Pi^{1}_{2}$ principles such as those above as problems, each having a class of instances, and for each instance, a class of solutions (see [16, Chapter 3] for a general discussion). Thus, for example, the instances of $\mathsf{RT}^{2}_{2}$ are all pairs $\langle X,c\rangle$ such that $X$ is an infinite set and $c:[X]^{2}\to 2$ is a coloring, with the solutions to $c$ being all infinite homogeneous sets $H\subseteq X$.
@@ 2. A proof of the Ginsburg–Sands theorem
Fix an infinite topological space $X$. By passing to a subspace if needed, we may assume $X$ is countable. For each $x\in X$, let $\operatorname{cl}(x)$ denote the closure of $x$ in $X$, i.e., the set of all $y\in X$ such that every open set containing $y$ also contains $x$. Then, for $x,y\in X$, define $x\sim y$ if $\operatorname{cl}(x)=\operatorname{cl}(y)$, which is an equivalence relation on $X$. If there is an $x$ such that the equivalence class $\{y\in\omega:x\sim y\}$ is infinite, then the subspace topology on this class is an indiscrete subspace of $X$. Moving forward, we can therefore assume that all equivalence classes under $\sim$ are finite. By choosing representatives, we may further assume that $\operatorname{cl}(x)\neq\operatorname{cl}(y)$ for all $x\neq y$ in $X$, so that $X$ itself is a $T_{0}$ space.
We now define a relation $\leq_{X}$ on $X$, as follows. For $x,y\in X$, set $x\leq_{X}y$ if $x\in\operatorname{cl}(y)$. Clearly, $\leq_{X}$ is reflexive and transitive, and our assumption that $X$ is $T_{0}$ implies that $\leq_{X}$ is antisymmetric. So, $\leq_{X}$ is a partial ordering. By $\mathsf{CAC}$, there exists an infinite subset of $X$ that either forms a chain under $\leq_{X}$ or an antichain under $\leq_{X}$. By regarding this as a subspace, we may assume this subset is in fact of all of $X$ for simplicity.
@@
First, suppose $X$ is a chain under $\leq_{X}$. Applying $\mathsf{ADS}$, we can fix an infinite subset $Y$ of $X$ that is either an ascending or descending sequence under $\leq_{X}$. Say $Y$ is an ascending sequence under $\leq_{X}$, so that its elements may be listed as $x_{0}<_{X}x_{1}<_{X}\cdots$. We claim that $Y$ under the subspace topology is homeomorphic to the final segment topology on $\omega$ via the map $x_{n}\mapsto n$. To see this, note first that if $n<m$ then every open set in $X$ containing $x_{n}$ also contains $x_{m}$ since $x_{n}\in\operatorname{cl}(x_{m})$. On the other hand, for every $n$ there is an open set containing $x_{n}$ but not any $x_{m}$ for $m<n$, since $x_{n}\notin\operatorname{cl}(x_{m})$ for any such $m$. It follows that the open sets in $Y$ are precisely the sets of the form $\{x_{m}:m\geq n\}$ for $n\in\omega$, which proves the claim. The case where $Y$ is a descending sequence under $\leq_{X}$ is symmetric, and produces a subspace homeomorphic to the initial segment topology.
We are thus left with the case where $X$ is an antichain. By definition of $\leq_{X}$, this means that $X$ is a $T_{1}$ space. Assume $X$ has no infinite subspace with the cofinite topology; we show it has an infinite discrete subspace. Since $X$ is $T_{1}$, every finite subset of $X$ is closed, and so every cofinite subset is open. Our assumption therefore implies that if $W$ is any infinite subspace of $X$, there exists an open set $U$ such that $U\cap W\neq\emptyset$ and $W\smallsetminus U$ is infinite. Let $U_{0}$ be any such open set for $W=X$, and let $x_{0}$ be any element of $U_{0}$. Next, suppose inductively that, for some $n\in\omega$, we have defined $x_{0},\ldots,x_{n}\in X$, along with open sets $U_{0},\ldots,U_{n}$ satisfying the following: $X\smallsetminus\bigcup_{i\leq n}U_{i}$ is infinite, and for all $i,j\leq n$, we have that $x_{i}\in U_{j}$ if and only if $i=j$. Letting $W=X\smallsetminus\bigcup_{i\leq n}U_{i}$, there exists an open set $U$ such that $U\cap W\neq\emptyset$ and $W\smallsetminus U$ is infinite. Let $x_{n+1}$ be any element of $U\cap W$, and let $U_{n+1}=U\smallsetminus\{x_{0},\ldots,x_{n}\}$. (Note that $U_{n+1}$ is open because $X$ is $T_{1}$.) Clearly, the inductive conditions are maintained. In this way, we thus define an infinite subspace $Y=\{x_{n}:n\in\omega\}$ of $X$, and this subspace is discrete since for each $n$ we have that $U_{n}\cap Y=\{x_{n}\}$ by construction. ∎